OpenStudy (lilsis76):

okay umm....how do i do this arithmetic sequence find the sum: pi/3 + 2pi/3 + pi + 4pi/3 +...+13pi/3 how do i do this? im so lost

OpenStudy (anonymous):

ignore the \(\pi\) and add \[\frac{1}{3}+\frac{2}{3}+\frac{3}{3}+...\] or better yet add \[1+2+3+...+13\] and the multiply the result by \(\frac{\pi}{3}\)

OpenStudy (lilsis76):

okay let me try it

OpenStudy (lilsis76):

haha im so lost hold on

OpenStudy (lilsis76):

91pi/3 ? @satellite73

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

yup you got it

OpenStudy (lilsis76):

coool :) now i got another problem on a diff proble m :?